### v2

@@ 2.2 Independent 3-D HOSVD
(5) $\mathcal{Y}\approx\mathcal{G}\times_{1}{X}_{1}\times_{2}{X}_{2}\times_{3}{X}_{3},$
where $\mathcal{G}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is the core tensor, and ${X}_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ is the $i$-th factor matrix such that ${X}_{i}^{\top}{X}_{i}=\mathbf{I}_{n_{i}}$. Note that $m_{i}\geq n_{i}$ and ${X}_{i}$ belongs to a Stiefel manifold, i.e., ${X}_{i}\in\mathbb{S}_{m_{i},n_{i}}$. By imposing orthogonality on the factor matrices, the decomposition in (5) can inherit some nice properties from the matrix SVD. For example, the core tensor can have the all-orthogonality and the ordering property [11].
@@
Figure 3: A stack of independent third order tensors
Also, a stack of independent matrices is denoted as $[X]\in\mathbb{R}^{m\times n\times N}$. And we say $[X]$ is independently orthogonal if $[{X}]^{\top}[{X}]=[\mathbf{I}_{n}]$, meaning $([{X}]^{(j)})^{\top}[{X}]^{(j)}=\mathbf{I}_{n}$, or equivalently $[X]\in[\mathbb{S}_{m,n}]$, meaning $[X]^{(j)}\in\mathbb{S}_{m,n}$. Then we define an independent (truncated) 3-D HOSVD of $[\mathcal{Y}]$ as
$[\mathcal{Y}]\approx[\mathcal{G}]\times_{1}[{X}_{1}]\times_{2}[{X}_{2}]\times_{3}[{X}_{3}],$
@@ 3 Proximal Block Coordinate Descent Algorithm
$\Phi(\mathcal{S},[{X}_{1}],[{X}_{2}],[{X}_{3}],[\mathcal{G}],\mathcal{L}):=$ $\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\varphi\left(\mathcal{S}\right)+\psi([\mathcal{G}])$
$+\frac{\delta}{2}H([X_{1}],[X_{2}],[X_{3}],[\mathcal{G}],\mathcal{L}),$
where $\varphi(\mathcal{S})=\|\mathcal{S}\|_{2,p}^{p}$, $\psi([\mathcal{G}])=\|[\mathcal{G}]\|_{1,w}$, and
(6) $H([X_{1}],[X_{2}],[X_{3}],[\mathcal{G}],\mathcal{L}):=\frac{1}{2}\|\mathbf{R}(\mathcal{L})-[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]\|^{2}_{F}.$
@@
$\mathcal{S}^{k+1}$ $\in\underset{\mathcal{S}}{\arg\!\min}\;\Phi(\mathcal{S},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\mathcal{S}}}{2}\|\mathcal{S}-\mathcal{S}^{k}\|^{2}_{F},$
$[{X}_{1}^{k+1}]$ $\in\underset{[{X}_{1}]\in[\mathbb{S}_{m_{1},n_{1}}]}{\arg\!\min}\!\Phi(\mathcal{S}^{k+1},[{X}_{1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{X}}{2}\|[{X}_{1}]-[{X}_{1}^{k}]\|^{2}_{F},$
$[{X}_{2}^{k+1}]$ $\in\underset{[{X}_{2}]\in[\mathbb{S}_{m_{2},n_{2}}]}{\arg\!\min}\!\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{X}}{2}\|[{X}_{2}]-[{X}_{2}^{k}]\|^{2}_{F},$
$[{X}_{3}^{k+1}]$ $\in\underset{[{X}_{3}]\in[\mathbb{S}_{m_{3},n_{3}}]}{\arg\!\min}\!\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{X}}{2}\|[{X}_{3}]-[{X}_{3}^{k}]\|^{2}_{F},$
$[\mathcal{G}^{k+1}]$ $=\underset{[\mathcal{G}]}{\arg\!\min}\;\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}],\mathcal{L}^{k})+\frac{\alpha_{\mathcal{G}}}{2}\|[\mathcal{G}]-[\mathcal{G}^{k}]\|^{2}_{F},$
$\mathcal{L}^{k+1}$ $=\underset{\mathcal{L}}{\arg\!\min}\;\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k+1}],\mathcal{L}),$
where parameters $\alpha_{\mathcal{S}},\alpha_{X},\alpha_{\mathcal{G}}>0$.
@@ 3.1 The update of $\mathcal{S}$
3.2 The update of $[X_{i}]$
Before we solve the optimization subproblem in terms of $[X_{i}]$ over independent Stiefel manifolds, we can rewrite its objective function using the following useful fact for unfolding of tensors
$\mathcal{Y}=\mathcal{G}\times_{i}X\mbox{ if and only if }{Y}_{(i)}=XG_{(i)}.$
Then by applying $X\in\mathbb{S}_{m,n}$, the Frobenious norm of tensors can be rewritten into the Frobenious norm of matrices
$\|\mathcal{G}\times_{i}X-\mathcal{L}\|_{F}^{2}=$ $\|XG_{(i)}-L_{(i)}\|_{F}^{2}$
(8) $=$ $\|G_{(i)}\|_{F}^{2}-2\langle XG_{(i)},L_{(i)}\rangle+\|L_{(i)}\|_{F}^{2},$
where $G_{(i)}$ and $L_{(i)}$ denote the mode-$i$ unfolding of $\mathcal{G}$ and $\mathcal{L}$, respectively. Since $\langle XG_{(i)},L_{(i)}\rangle=\langle X,L_{(i)}G_{(i)}^{\top}\rangle$ and $\|X\|_{F}^{2}=n$, minimizing $\|\mathcal{G}\times_{i}X-\mathcal{L}\|_{F}^{2}$ over $X$ on the Stiefel manifold is equivalent to minimizing $\|X-L_{(i)}G_{(i)}^{\top}\|_{F}^{2}$ over the Stiefel manifold. Hence, $[X_{i}^{k+1}]$ can be computed via the projection of unfolding matrices onto the Stiefel manifolds independently as follows
(9) $[X_{i}^{k+1}]\in\operatorname{Proj}_{[\mathbb{S}_{m_{i},n_{i}}]}\left([X_{i}^{k}]-\tilde{\alpha}_{X}\left([X_{i}^{k}]-[P_{i}^{k}][Q_{i}^{k}]^{\top}\right)\right),$
where $[P_{i}^{k}]=(\mathbf{R}(\mathcal{L}^{k}))_{[(i)]}$, $[Q_{1}^{k}]=\left([\mathcal{G}^{k}]\times_{2}[X^{k}_{2}]\times_{3}[X^{k}_{3}]\right)_{[(1)]}$, $[Q_{2}^{k}]=\big{(}[\mathcal{G}^{k}]\times_{1}[X_{1}^{k+1}]$$\times_{3}[X^{k}_{3}]\big{)}_{[(2)]}$, $[Q_{3}^{k}]=\left([\mathcal{G}^{k}]\times_{1}[X^{k+1}_{1}]\times_{2}[X^{k+1}_{2}]\right)_{[(3)]}$, and parameter $\tilde{\alpha}_{X}=\frac{\delta}{\delta+\alpha_{X}}$.
@@ 3.2 The update of $[X_{i}]$
Given $A\in\mathbb{R}^{m\times n}$, $m\geq n$, consider the following Stiefel manifold projection problem
(10) $\begin{split}&\min_{X\in\mathbb{R}^{m\times n}}\|X-A\|^{2}_{F}\\ &{\rm s.t.}\quad X^{\top}X=\mathbf{I}_{n}.\end{split}$
Then the set of optimal solutions of problem (10), denoted as $\Omega^{*}(A)$, is given by
$\Omega^{*}(A)=\{UV^{\top}|$ $A=U\Sigma V^{\top},U\in\mathbb{R}^{m\times n},\Sigma\in\mathbb{R}^{n\times n},V\in\mathbb{R}^{n\times n}$
@@
where $U\Sigma V^{\top}$ is a reduced SVD of $A$ and $\sigma(A)\in\mathbb{R}^{n}$ is a vector of all the singular values of $A$. In particular, if $A$ is of full column rank $n$, then $\Omega^{*}(A)$ is a singleton.
According to Lemma 3.1, problem (10) has a closed form solution, even though it may have multiple solutions when the given matrix does not have full column rank. Hence, if $[U][\Sigma][V]^{\top}$ is an independent reduced SVD of $[A]\in\mathbb{R}^{m\times n\times N}$, then $[U][V]^{\top}\in\operatorname{Proj}_{[\mathbb{S}_{m,n}]}([A])$.
@@ 3.3 The update of $[\mathcal{G}]$
The subproblem for updating $[\mathcal{G}]$ can be reformulated by using the following property that for any $X\in\mathbb{S}_{m,n}$, $m\geq n$,
$\|\mathcal{G}\times_{i}X-\mathcal{L}\|_{F}^{2}$
$=$ $\|G_{(i)}-X^{\top}L_{(i)}\|_{F}^{2}-\|X^{\top}L_{(i)}\|_{F}^{2}+\|L_{(i)}\|_{F}^{2}$
$=$ $\|\mathcal{G}-\mathcal{L}\times_{i}X^{\top}\|_{F}^{2}-\|\mathcal{L}\times_{i}X^{\top}\|_{F}^{2}+\|\mathcal{L}\|_{F}^{2},$
which is derived from (8) and the constraint that $X^{\top}X=\mathbf{I}_{n}$. Then $[\mathcal{G}^{k+1}]$ can be computed by
(11) $[\mathcal{G}^{k+1}]=\operatorname{prox}_{\tilde{w}_{\mathcal{G}}\psi}\left([\mathcal{G}^{k}]-\tilde{\alpha}_{\mathcal{G}}([\mathcal{G}^{k}]-[\mathcal{O}^{k+1}])\right),$
where $[\mathcal{O}^{k+1}]=\mathbf{R}(\mathcal{L}^{k})\times_{1}[X_{1}^{k+1}]^{\top}\times_{2}[X_{2}^{k+1}]^{\top}\times_{3}[X_{3}^{k+1}]^{\top}$, $\tilde{w}_{\mathcal{G}}=\frac{1}{\delta+\alpha_{\mathcal{G}}}$, and $\tilde{\alpha}_{\mathcal{G}}=\frac{\delta}{\delta+\alpha_{\mathcal{G}}}$.

### v3

@@ 2.2 Independent 3-D HOSVD
(5) $\mathcal{Y}\approx\mathcal{G}\times_{1}\boldsymbol{X}_{1}\times_{2}\boldsymbol{X}_{2}\times_{3}\boldsymbol{X}_{3},$
where $\mathcal{G}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is the core tensor, and $\boldsymbol{X}_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ is the $i$-th factor matrix such that $\boldsymbol{X}_{i}^{\top}\boldsymbol{X}_{i}=\mathbf{I}_{n_{i}}$. Note that $m_{i}\geq n_{i}$ and $\boldsymbol{X}_{i}$ belongs to a Stiefel manifold, i.e., $\boldsymbol{X}_{i}\in\mathbb{S}_{m_{i},n_{i}}$. By imposing orthogonality on the factor matrices, the decomposition in (5) can inherit some nice properties from the matrix SVD. For example, the core tensor can have the all-orthogonality and the ordering property [11].
@@
Figure 3: A stack of independent tensors.
Next, we say $[\boldsymbol{X}]$ is independently orthogonal if $[\boldsymbol{X}]^{\top}[\boldsymbol{X}]=[\mathbf{I}_{n}],\;$ meaning $([\boldsymbol{X}]^{(j)})^{\top}[\boldsymbol{X}]^{(j)}=\mathbf{I}_{n}$, or equivalently $[\boldsymbol{X}]\in[\mathbb{S}_{m,n}]$, meaning $[\boldsymbol{X}]^{(j)}\in\mathbb{S}_{m,n}$. Then we define an independent (truncated) 3-D HOSVD of $[\mathcal{Y}]$ as
(6) $[\mathcal{Y}]\approx[\mathcal{G}]\times_{1}[\boldsymbol{X}_{1}]\times_{2}[\boldsymbol{X}_{2}]\times_{3}[\boldsymbol{X}_{3}],$
@@
The core tensor $\mathcal{G}$ in the 3-D HOSVD in (5), as well as the stack of independent core tensors $[\mathcal{G}]$ in (6), play important roles in characterizing the low-rankness of the tensors $\mathcal{Y}$ and $[\mathcal{Y}]$, respectively. Specifically, the core tensors and their entries reflect the level of interaction among the different components [22]. Therefore, low-rankness can be promoted by imposing the sparsity on the core tensors.
@@ 3 Proximal Block Coordinate Descent Algorithm
$\Phi(\mathcal{S},[\boldsymbol{X}_{1}],[\boldsymbol{X}_{2}],[\boldsymbol{X}_{3}],[\mathcal{G}],\mathcal{L}):=$ $\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\varphi\left(\mathcal{S}\right)+\psi([\mathcal{G}])$
$+\delta H([\boldsymbol{X}_{1}],[\boldsymbol{X}_{2}],[\boldsymbol{X}_{3}],[\mathcal{G}],\mathcal{L}),$
where $\varphi(\mathcal{S})=\|\mathcal{S}\|_{2,p}^{p}$, $\psi([\mathcal{G}])=\|[\mathcal{G}]\|_{1,\boldsymbol{w}}$, and
(7) $H([\boldsymbol{X}_{1}],[\boldsymbol{X}_{2}],[\boldsymbol{X}_{3}],[\mathcal{G}],\mathcal{L}):=\frac{1}{2}\|\mathscr{R}(\mathcal{L})-[\mathcal{G}]\times_{1}[\boldsymbol{X}_{1}]\times_{2}[\boldsymbol{X}_{2}]\times_{3}[\boldsymbol{X}_{3}]\|^{2}_{F}.$
@@
$\mathcal{S}^{k+1}$ $\in\underset{\mathcal{S}}{\arg\!\min}\;\Phi(\mathcal{S},[\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\mathcal{S}}}{2}\|\mathcal{S}-\mathcal{S}^{k}\|^{2}_{F},$
$[\boldsymbol{X}_{1}^{k+1}]$ $\in\underset{[\boldsymbol{X}_{1}]\in[\mathbb{S}_{m_{1},n_{1}}]}{\arg\!\min}\!\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\boldsymbol{X}}}{2}\|[\boldsymbol{X}_{1}]-[\boldsymbol{X}_{1}^{k}]\|^{2}_{F},$
$[\boldsymbol{X}_{2}^{k+1}]$ $\in\underset{[\boldsymbol{X}_{2}]\in[\mathbb{S}_{m_{2},n_{2}}]}{\arg\!\min}\!\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\boldsymbol{X}}}{2}\|[\boldsymbol{X}_{2}]-[\boldsymbol{X}_{2}^{k}]\|^{2}_{F},$
$[\boldsymbol{X}_{3}^{k+1}]$ $\in\underset{[\boldsymbol{X}_{3}]\in[\mathbb{S}_{m_{3},n_{3}}]}{\arg\!\min}\!\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\boldsymbol{X}}}{2}\|[\boldsymbol{X}_{3}]-[\boldsymbol{X}_{3}^{k}]\|^{2}_{F},$
$[\mathcal{G}^{k+1}]$ $=\underset{[\mathcal{G}]}{\arg\!\min}\;\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}],\mathcal{L}^{k})+\frac{\alpha_{\mathcal{G}}}{2}\|[\mathcal{G}]-[\mathcal{G}^{k}]\|^{2}_{F},$
$\mathcal{L}^{k+1}$ $=\underset{\mathcal{L}}{\arg\!\min}\;\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k+1}],\mathcal{L}),$
where parameters $\alpha_{\mathcal{S}},\alpha_{\boldsymbol{X}},\alpha_{\mathcal{G}}>0$.
@@ 3.1 The update of $\mathcal{S}$
3.2 The update of $[\boldsymbol{X}_{i}]$
Before we solve the optimization subproblem in terms of $[\boldsymbol{X}_{i}]$ over independent Stiefel manifolds, we can rewrite its objective function using the following useful fact for unfolding of tensors
$\mathcal{Y}=\mathcal{G}\times_{i}\boldsymbol{X}\mbox{ if and only if }\boldsymbol{Y}_{(i)}=\boldsymbol{X}\boldsymbol{G}_{(i)}.$
Then by applying $\boldsymbol{X}\in\mathbb{S}_{m,n}$, the Frobenious norm of tensors can be rewritten into the Frobenious norm of matrices
$\|\mathcal{G}\times_{i}\boldsymbol{X}-\mathcal{L}\|_{F}^{2}=$ $\|\boldsymbol{X}\boldsymbol{G}_{(i)}-\boldsymbol{L}_{(i)}\|_{F}^{2}$
(9) $=$ $\|\boldsymbol{G}_{(i)}\|_{F}^{2}-2\langle\boldsymbol{X}\boldsymbol{G}_{(i)},\boldsymbol{L}_{(i)}\rangle+\|\boldsymbol{L}_{(i)}\|_{F}^{2},$
where $\boldsymbol{G}_{(i)}$ and $\boldsymbol{L}_{(i)}$ denote the mode-$i$ unfolding of $\mathcal{G}$ and $\mathcal{L}$, respectively. Since $\langle\boldsymbol{X}\boldsymbol{G}_{(i)},\boldsymbol{L}_{(i)}\rangle=\langle\boldsymbol{X},\boldsymbol{L}_{(i)}\boldsymbol{G}_{(i)}^{\top}\rangle$ and $\|\boldsymbol{X}\|_{F}^{2}=n$, minimizing $\|\mathcal{G}\times_{i}\boldsymbol{X}-\mathcal{L}\|_{F}^{2}$ over $X$ on the Stiefel manifold is equivalent to minimizing $\|\boldsymbol{X}-\boldsymbol{L}_{(i)}\boldsymbol{G}_{(i)}^{\top}\|_{F}^{2}$ over the Stiefel manifold. Hence, $[\boldsymbol{X}_{i}^{k+1}]$ can be computed via the projection of unfolding matrices onto the Stiefel manifolds independently as follows
(10) $[\boldsymbol{X}_{i}^{k+1}]\in\operatorname{Proj}_{[\mathbb{S}_{m_{i},n_{i}}]}\left([\boldsymbol{X}_{i}^{k}]-\tilde{\alpha}_{\boldsymbol{X}}\left([\boldsymbol{X}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)\right),$
where $[\boldsymbol{P}_{i}^{k}]=(\mathscr{R}(\mathcal{L}^{k}))_{[(i)]}$, $[\boldsymbol{Q}_{1}^{k}]=\left([\mathcal{G}^{k}]\times_{2}[\boldsymbol{X}^{k}_{2}]\times_{3}[\boldsymbol{X}^{k}_{3}]\right)_{[(1)]}$, $[\boldsymbol{Q}_{2}^{k}]=\big{(}[\mathcal{G}^{k}]\times_{1}[\boldsymbol{X}_{1}^{k+1}]$$\times_{3}[\boldsymbol{X}^{k}_{3}]\big{)}_{[(2)]}$, $[\boldsymbol{Q}_{3}^{k}]=\left([\mathcal{G}^{k}]\times_{1}[\boldsymbol{X}^{k+1}_{1}]\times_{2}[\boldsymbol{X}^{k+1}_{2}]\right)_{[(3)]}$, and parameter $\tilde{\alpha}_{\boldsymbol{X}}=\frac{\delta}{\delta+\alpha_{\boldsymbol{X}}}$.
@@ 3.2 The update of $[\boldsymbol{X}_{i}]$
Given $\boldsymbol{A}\in\mathbb{R}^{m\times n}$, $m\geq n$, consider the following Stiefel manifold projection problem
(11) $\begin{split}&\min_{\boldsymbol{X}\in\mathbb{R}^{m\times n}}\|\boldsymbol{X}-\boldsymbol{A}\|^{2}_{F}\\ &{\rm s.t.}\quad\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{n}.\end{split}$
Then the set of optimal solutions of problem (11), denoted as $\Omega^{*}(\boldsymbol{A})$, is given by
$\Omega^{*}(\boldsymbol{A})=\{\boldsymbol{U}\boldsymbol{V}^{\top}|$ $\boldsymbol{A}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{\top},\boldsymbol{U}\in\mathbb{R}^{m\times n},\boldsymbol{\Sigma}\in\mathbb{R}^{n\times n},\boldsymbol{V}\in\mathbb{R}^{n\times n}$
@@
where $\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{\top}$ is a reduced SVD of $\boldsymbol{A}$ and $\boldsymbol{\sigma}(\boldsymbol{A})\in\mathbb{R}^{n}$ is a vector of all the singular values of $\boldsymbol{A}$. In particular, if $\boldsymbol{A}$ is of full column rank $n$, then $\Omega^{*}(\boldsymbol{A})$ is a singleton.
According to Lemma 3.1, problem (11) has a closed form solution, even though it may have multiple solutions when the given matrix does not have full column rank. Hence, if $[\boldsymbol{U}][\boldsymbol{\Sigma}][\boldsymbol{V}]^{\top}$ is an independent reduced SVD of $[\boldsymbol{A}]\in\mathbb{R}^{m\times n\times N}$, then $[\boldsymbol{U}][\boldsymbol{V}]^{\top}\in\operatorname{Proj}_{[\mathbb{S}_{m,n}]}([\boldsymbol{A}])$.
@@ 3.3 The update of $[\mathcal{G}]$
The subproblem for updating $[\mathcal{G}]$ can be reformulated by using the following property that for any $\boldsymbol{X}\in\mathbb{S}_{m,n}$, $m\geq n$,
$\|\mathcal{G}\times_{i}\boldsymbol{X}-\mathcal{L}\|_{F}^{2}$
$=$ $\|\boldsymbol{G}_{(i)}-\boldsymbol{X}^{\top}L_{(i)}\|_{F}^{2}-\|\boldsymbol{X}^{\top}\boldsymbol{L}_{(i)}\|_{F}^{2}+\|\boldsymbol{L}_{(i)}\|_{F}^{2}$
$=$ $\|\mathcal{G}-\mathcal{L}\times_{i}\boldsymbol{X}^{\top}\|_{F}^{2}-\|\mathcal{L}\times_{i}\boldsymbol{X}^{\top}\|_{F}^{2}+\|\mathcal{L}\|_{F}^{2},$
which is derived from (9) and the constraint that $\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{n}$. Then $[\mathcal{G}^{k+1}]$ can be computed by
$[\mathcal{G}^{k+1}]=$ $\underset{[\mathcal{G}]}{\arg\!\min}\;\psi([\mathcal{G}])+\frac{\delta}{2}\|[\mathcal{G}]-[\mathcal{O}^{k+1}]\|^{2}_{F}+\frac{\alpha_{\mathcal{G}}}{2}\|[\mathcal{G}]-[\mathcal{G}^{k}]\|^{2}_{F},$
$=$ $\underset{[\mathcal{G}]}{\arg\!\min}\;\tilde{w}_{\mathcal{G}}\psi([\mathcal{G}])+\frac{1}{2}\|[\mathcal{G}]-\left([\mathcal{G}^{k}]-\tilde{\alpha}_{\mathcal{G}}([\mathcal{G}^{k}]-[\mathcal{O}^{k+1}])\right)\|^{2}_{F},$
(12) $=$ $\operatorname{prox}_{\tilde{w}_{\mathcal{G}}\psi}\left([\mathcal{G}^{k}]-\tilde{\alpha}_{\mathcal{G}}([\mathcal{G}^{k}]-[\mathcal{O}^{k+1}])\right),$
where $[\mathcal{O}^{k+1}]=\mathscr{R}(\mathcal{L}^{k})\times_{1}[\boldsymbol{X}_{1}^{k+1}]^{\top}\times_{2}[\boldsymbol{X}_{2}^{k+1}]^{\top}\times_{3}[\boldsymbol{X}_{3}^{k+1}]^{\top}$, $\tilde{w}_{\mathcal{G}}=\frac{1}{\delta+\alpha_{\mathcal{G}}}$, and $\tilde{\alpha}_{\mathcal{G}}=\frac{\delta}{\delta+\alpha_{\mathcal{G}}}$.